Solve the equation 3^2x = 7(3^x)- 12. write final solution as exact or in form of x= log(base a)b
*Would I take the logs of both sides and then use log properties?

Answer by lwsshak3(11628)

Solve the equation 3^2x = 7(3^x)- 12
**
7(3^x)-3^2x=12
factor out 3^x
(3^x)(7-3^x)=12
take log of both sides:
log(3^x)+log(7-3^x)=log(12)=1.0792
log[(3^x)(7-3^x)]=1.0792
convert to exponential form:
10^1.0792=(3^x)(7-3^x)=7(3^x)-3^2x=12
7(3^x)-3^2x-12=0
let u=3^x
u^2=3^2x
u^2-7u+12=0
(u-4)(u+3)=0
u=-3≠3^x (reject)
or
u=4=3^x
xlog3=log4
x=log4/log3≈1.2619.. (ans)